OpenStudy (anonymous):

E. Evaluate the given expression. Show work for partial credit. (5 points) a = -4, b = 8, c = 2, and d = -3 |ad-c|/|b-3|

OpenStudy (anonymous):

The answer is 2

OpenStudy (anonymous):

ok but can you help me work it out

OpenStudy (anonymous):

@bhhoward

OpenStudy (anonymous):

[-4*-3 - 2] / [ 8 - 3 ] -4 * -3 = 12 - 2 = 10 8-3=5 10 / 5 = 2

OpenStudy (anonymous):

Thankyou so much

OpenStudy (anonymous):

No problem :)